OpenStudy (1018):

y=1-2u u = x^3 - 2x^2 + 6 find y''(x=1)

OpenStudy (welshfella):

y = 1 - 2( x^3 - 2x^2 + 6) y = 1 - 2x^3 + 4x^2 - 12 what did you get for y' ?

OpenStudy (1018):

hey sorry late. are you still there? i tried to get y'' with that one that the answer isnt one of the choices